OpenStudy (anonymous):

Help for Metal Please Consider the following pair of equations: −x − y = −5 y = x + 1 If the two equations are graphed, at what point do the lines representing the two equations intersect? (−2, 3) (3, −2) (2, 3) (3, 2)

iYuko (iyuko):

I don't want metal.

OpenStudy (anonymous):

ok

OpenStudy (haseeb96):

There are two equations in this question we call -x -y =-5 x+y=5 ...........eq#1 since,- signs are cancelled from each side y=x+1............eq#2 now put the value of y of eq#2 into eq#1 at the place y it would be x +x+1 =5 2x = 5-1 2x=4 x=2 now put x=2 in eq#2 you will get the value of y y= 2+1 y=3 so answer will be (x,y)=(2,3)

OpenStudy (haseeb96):

If you have any question, ask me

OpenStudy (haseeb96):

@TheStudent2001

OpenStudy (anonymous):

Well I am not the best at prealgebra so I need help with this

OpenStudy (anonymous):

-y-x =-5 (1) multiply by minus to get: -(-y) = -(x-5) which becomes y + x= 5 so you have -y = x -5 (1) and y=x +1 (2) Now to things that are equal to the same thing are equal to each other so: x-5 =x+1 can you solve that?

OpenStudy (anonymous):

ok I understand it so far but I dont understand what else to do

OpenStudy (anonymous):

bring x's to one side and integers to the other side

OpenStudy (anonymous):

o

OpenStudy (anonymous):

Is that o good! or o bad?

OpenStudy (anonymous):

Good

OpenStudy (lynfran):

hey do you want to know a faster trick to complete this...like...just plug in your answer option into the equation... -x-y= -5 ......-(2)-3=? y=x+1 .... y=(2)+3...

OpenStudy (anonymous):

Ok I will try it

OpenStudy (anonymous):

Thank You for yalls help

OpenStudy (anonymous):

yes and here is an even faster way: -y-x = -5 y=x+1 ------------ -y-x = -5 y-x = 1 --------- Add them and the y's cancel, leaving -2x = -4 =>x=2